OpenStudy (anonymous):

Simplify the following: (x – 5)/(x2 – x – 20). A. x - 4 B. x + 4 C. x2 - 15 D. 1/(x + 4)

OpenStudy (anonymous):

(x - 5)/(x^2 - x - 20) = (x - 5)/(x^2 + 4x - 5x - 20) = (x - 5)/[(x^2 + 4x) - (5x + 20)] = (x - 5)/[x(x + 4) - 5(x + 4)] = (x - 5)/(x + 4)(x - 5) (cancel out x - 5) = 1/(x + 4)

OpenStudy (anonymous):

So D

OpenStudy (anonymous):

Obv lol

OpenStudy (anonymous):

Dont get huit

OpenStudy (anonymous):

wow you have me Scared..

OpenStudy (anonymous):

Better be lololo jkjkjk

OpenStudy (anonymous):

gimmi a Freaking medal ( ;

OpenStudy (anonymous):

ill taske it away be good